40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1909     Entry time: Sat Aug 15 05:08:55 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Friday night locking 
Summary: DD hand off fails for DRFPMI.

Tonight, I did a lot of house keeping work.

(1) I noticed that the reference cavity (RC) was locked to TEM10.
This was probably the reason why we had to increase the FSS common gain.
I re-locked the RC to TEM00. Now the common gain value is back to the original.

(2) The MC WFS did not engage. I found that c1dcuepics had the /cvs/cds mounting problem.
I rebooted it. Then MC WFS started working.

(3) After checking that the MC WFS QPDs are centered for direct reflection (the MZ half fringe method),
I locked the MC and tweaked the mirror alignment (mainly MC3) to offload the WFS feedback signals.
Now the MC locks to TEM00 robustly.

(4) Since the MC mirror alignment is touchy recently, I did not like the idea of mis-aligning MC2
when you do the LSC PD offset adjustment. So I modified the LSCoffset script so that it will close
the PSL shutter instead of mis-aligning MC2.

(5) I changed the PD11_Q criteria for success in the alignment scripts because PD11_Q is now lower
than before due to the lower laser power.

(6) Since today's bootfest, some epics values were not properly restored. Some of the PD gains were
unmatched between I and Q. I corrected these with the help of conlog.

(7) By checking the open loop TFs, I found that the short DOFs have significantly lower UGFs than before,
probably due to the lower laser power. I increased the gains of MICH, PRCL and SRCL by a factor of 2 for
the full configuration.
For the DRM configuration the changes I made were:

PRC -0.15 -> -0.3
SRC 0.2 -> 0.6
MICH 0.5 -> 0.5

(8) I locked the DRFPMI with arm offsets, then adjusted the demodulation phases of PD6,PD7,PD8 and PD9 (DD PDs)
to minimize the offsets in the error signal, while locked with the single demodulation signals.

Change log:
PD6_PHASE 201 -> 270
PD7_PHASE 120 -> 105
PD8_PHASE 131 -> 145
PD9_PHASE -45 -> -65


(9) I ran senseDRM to get the sensing Matrix for the short DOFs using DD signals in DRM configuration.

(10) Still the DD hand off fails for DRFPMI. It succeeds for DRM.
ELOG V3.1.3-